# Diffusion-Based Impedance Learning for Contact-Rich Manipulation Tasks

### v2

@@ V Discussion
The experimental results confirm the central goal of this work: enabling robust contact-rich manipulation by bridging the information and energy domains and unifying state-of-the-art generative AI algorithms with model-based control. In particular, the success of peg-in-hole insertions with unseen geometries demonstrates that sZFT reconstruction generalizes equilibrium adaptation beyond the training distribution.
Compared to prior learning-based approaches, our framework explicitly incorporates energy consistency, enabling stable impedance adaptation even under contact. Conversely, compared to traditional model-based Impedance Controllers, it gains adaptability by reconstructing equilibrium trajectories directly from data. This ability to bridge domains accounts for the strong performance observed in unstructured environments.
A key element of the framework is directional stiffness adaptation. This mechanism is generic and can be tailored to applications in other domains. One example is robotic-assisted upper-limb therapy, where stiffness must remain high along non-movement directions to ensure patient safety, while stiffness along the intended movement direction should adapt depending on whether the patient can execute the motion independently. In such settings, our framework could provide both safety and flexibility, adjusting assistance in real time.
@@
The proposed approach guarantees passivity when stiffness is reduced: lowering stiffness dissipates stored potential energy and ensures that interaction with a passive environment remains output strictly passive. However, when stiffness is increased back toward the baseline, passivity is not generally guaranteed, as such changes can inject additional energy whenever displacements are nonzero. In these cases, the controller can be combined with our earlier work on passive impedance shaping [61] to enforce strict passivity.
Overall, our findings demonstrate that Diffusion-Based Impedance Learning is highly effective in the experimental setups investigated. While this provides strong evidence for the viability of the approach, further studies across a wider range of tasks and environments are needed to fully establish its generality. Nevertheless, the framework offers a promising design principle for stable, safe, and adaptable robot interaction in other domains like industrial assembly, physical Human-Robot Interaction, and robotic-assisted therapy.
@@ VI Conclusion
The framework consistently outperformed fixed and heuristic strategies in experiments on a KUKA LBR iiwa. It enabled smooth traversal of obstacle-rich parkour environments where conventional Impedance Control failed, and achieved 30/30 success rates in cylindrical, square, and star peg-in-hole insertions without any task-specific demonstrations.
Beyond these results, Diffusion-Based Impedance Learning establishes a general interface between learning and model-based control. It extends the role of generative models from visuomotor trajectory planning to physical interaction.
Looking ahead, we aim to extend this framework from single-arm to dual-arm setups and to human–robot collaborative systems. Embedding Diffusion-Based Impedance Learning in hierarchical architectures that couple high-level visuomotor policies with low-level impedance regulation could yield robots that both plan intelligently and interact safely in unstructured, human-centered environments. Ultimately, such systems may handle the variability of industrial assembly lines, support cooperative human–robot tasks, and provide personalized assistance in clinical care.
Acknowledgements
We gratefully acknowledge the support of KUKA, an international leader in automation solutions, and specifically thank them for providing the KUKA robots used in our experiments.
Appendix A Energy-based Impedance Controller
A-A Elastic Force
The translational restoring force arises $\bm{f}\in\mathbb{R}^{3}$ from the displacement between the end-effector position $\bm{p}\in\mathbb{R}^{3}$ and the reference ZFT position $\bm{p}_{0}\in\mathbb{R}^{3}$. It is modeled as
$\bm{f}=\bm{K}_{t}\ (\bm{p}_{0}-\bm{p})-\bm{B}_{t}\ \dot{\bm{p}},$ (29)
where $\bm{K}_{t}\in\mathbb{R}^{3\times 3}$ is the translational stiffness matrix, $\bm{B}_{t}\in\mathbb{R}^{3\times 3}$ the damping matrix, and $\dot{\bm{p}}\in\mathbb{R}^{3}$ the end-effector translational velocity.
For implementation, the damping is expressed proportional to stiffness as
$\bm{B}_{t}=\lambda_{t}\,\bm{K}_{t},$ (30)
with $\lambda_{t}$ the translational positive time constant.
A-B Elastic Moment
Rotational restoring moments $\bm{m}\in\mathbb{R}^{3}$ are expressed relative to the ZFT orientation $\mathbf{q}_{0}\in SU(2)$, represented in axis–angle form as $(\bm{u}_{0},\theta_{0})$. The elastic moment is given by
$\bm{m}=\bm{K}_{r}\,\theta_{0}\bm{u}_{0}-\bm{B}_{r}\ \bm{\omega},$ (31)
where $\bm{K}_{r}\in\mathbb{R}^{3\times 3}$ is the rotational stiffness matrix, $\bm{B}_{r}\in\mathbb{R}^{3\times 3}$ the rotational damping matrix, and $\bm{\omega}\in\mathbb{R}^{3}$ the angular velocity of the end-effector.
Analogous to translations, the damping term is chosen proportional to stiffness:
$\bm{B}_{r}=\lambda_{r}\,\bm{K}_{r},$ (32)
with $\lambda_{r}$ the rotational positive time constant.
A-C Damping design
The damping design was formulated to ensure that energy dissipation scales consistently with both the commanded stiffness $\bm{K}\in\mathbb{R}^{3\times 3}$ and the task-space inertia $\bm{\Lambda}\in\mathbb{R}^{3\times 3}$. This construction was applied in the translational and rotational subspaces alike. To account for numerical issues near singular configurations, $\bm{\Lambda}$ was computed using a damped least-squares method.
$\bm{D}=d\,\bm{I}_{3}.$ (33)
The symmetric positive semidefinite square roots $\sqrt{\bm{\Lambda}}$ and $\sqrt{\bm{K}}$ are obtained by eigendecomposition for $\bm{\Lambda}$ and elementwise for $\bm{K}$. As presented in [62], an intermediate damping matrix is constructed as
$\bm{b}\;=\;\sqrt{\bm{\Lambda}}\,\bm{D}\,\sqrt{\bm{K}}\;+\;\sqrt{\bm{K}}\,\bm{D}\,\sqrt{\bm{\Lambda}}.$ (34)
From $\bm{b}$, a scalar damping coefficient is computed to normalize the dissipation relative to the overall stiffness:
$\lambda\;=\;\frac{2\,\mathrm{tr}(\bm{b})}{\mathrm{tr}(\bm{K})}\;=\;\frac{2\,\mathrm{tr}(\bm{b})}{k_{x}+k_{y}+k_{z}}.$ (35)
The final damping matrix can then be expressed as
$\bm{B}\;=\;\lambda\,\bm{K}.$ (36)
Appendix B ZFT-based Telemanipulation with Apple Vision Pro
@@
Figure 10: Pose tracking and visualization with Apple Vision Pro. Left: operator performing telemanipulation. Right: virtual reconstruction of wrist, hand, and head poses in a 3D scene. [27]
@@
Pose data were transferred via shared memory between the Python-based VisionProTeleop module and the C++ robot controller. In this way, the Apple Vision Pro provided a seamless method to generate sZFT ground truth data through markerless and physically unconstrained teleoperation. The implementation of the teleoperation interface is available in our GitHub repository [^11] https://github.com/StrokeAIRobotics/DiffusionBasedImpedanceLearning.
Appendix C Peg-In-Hole experiments
This appendix complements the main results by reporting stiffness adaptation for the circular and square pegs. The corresponding results for the star peg are already presented in the main text (Fig. 9). The experimental setup, controllers, and parameters matched the main text: baseline stiffness $K_{t,i,\text{max}}=$800\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$\$ and $K_{r,i,\text{max}}=$150\text{\,}\mathrm{N}\mathrm{m}\mathrm{/}\mathrm{r}\mathrm{a}\mathrm{d}$\$, damping as in Appendix A-C, and force/moment thresholds of $1\text{\,}\mathrm{N}$ and $1\text{\,}\mathrm{N}\mathrm{m}$. All quantities are expressed in the robot base frame. The lower plots in both of the following figures show the relative translational and rotational factors $\rho_{t,i}$ and $\rho_{r,i}$ from (27), which modulate the reduction terms in the adaptive law.
@@ Appendix C Peg-In-Hole experiments
Figure 11: Insertion of a circular peg. Measured external wrenches and commanded translational and rotational stiffness during insertion. The lower plots show the relative translational and rotational factors (Eq. (27)) and their influence on stiffness modulation. Rotational adjustments are small due to axial symmetry; stiffness reductions occur primarily along directions with low task relevance.
Figure 12: Insertion of a quare peg. Measured external wrenches and commanded translational and rotational stiffness during insertion. The lower plots show the relative translational and rotational factors (Eq. (27)). Directional reductions are more pronounced than for the circular peg, reflecting the need to relieve off-axis contact while preserving stiffness along the insertion and alignment axes.

### v3

@@ V Discussion
The experimental results confirm the central goal of this work: enabling robust contact-rich manipulation by bridging the information and the energy domains and therefore unifying state-of-the-art generative AI algorithms with model-based control. In particular, the success of peg-in-hole insertions with unseen tool and task geometries demonstrates that sZFT reconstruction generalizes beyond the training distribution.
Compared to prior learning-based approaches, our framework explicitly incorporates energy transformation, enabling stable impedance adaptation under contact. Compared to traditional model-based impedance controllers, it gains adaptability by reconstructing equilibrium trajectories directly from data. This ability to bridge domains accounts for the strong performance observed in unstructured environments.
A key element of the framework is directional stiffness adaptation. This mechanism is generic and can be tailored to applications in other domains. One example is robotic-assisted upper-limb therapy, where stiffness must remain high along non-movement directions, while stiffness along the intended movement direction should adapt depending on whether the patient can execute the motion independently. In such settings, the same framework can be used by solely adapting the directional stiffness adaptation.
The force and moment thresholds in the stiffness estimator affect the responsiveness of impedance adaptation. In our experiments, values of 1 $\mathrm{N}$ and 1 $\mathrm{N}\mathrm{m}$ were selected to match the noise characteristics of the ATI Gamma force/torque sensor. These thresholds are not task-specific and can be adjusted for different sensors or applications.
@@
The proposed approach guarantees passivity when stiffness is reduced: lowering stiffness dissipates stored potential energy and ensures that interaction with a passive environment remains output strictly passive. However, when stiffness is increased back toward the baseline, passivity is not generally guaranteed, as such changes can inject additional energy whenever displacements are nonzero. In these cases, the controller can be combined with previous work on passive impedance shaping [34] to enforce strict passivity.
In our work, we adopt a Transformer-based Diffusion Model due to its ability to capture long range temporal dependencies and flexibly condition on multimodal interaction signals. Future work should investigate a systematic comparison with alternative architectures such as convolutional or hybrid models under matched parameter and computational budgets.
Overall, our findings demonstrate that Diffusion-Based Impedance Learning is highly effective in the experimental setups investigated. While this provides strong evidence for the viability of the approach, further studies across a wider range of tasks and environments are needed to fully establish its generality. Nevertheless, the framework offers a promising design principle for stable, safe, and adaptable robot interaction in other domains inclduing industrial assembly, physical Human-Robot Interaction, and robotic-assisted rehabilitation.
@@ VI Conclusion
The framework consistently outperformed fixed and heuristic strategies in experiments on a KUKA LBR iiwa. It enabled smooth traversal of obstacle-rich parkour environments where conventional impedance control failed, and achieved 30/30 success rates in cylindrical, square, and star peg-in-hole insertions without any task-specific demonstrations in the training data. The experiments in this work intentionally withhold geometric information from the controller, targeting scenarios in which such knowledge is unavailable or unreliable due to unstructured environments. Diffusion-Based Impedance Learning should therefore be understood as a general, interaction-driven alternative rather than a replacement for highly engineered solutions in fully structured assembly lines.
Beyond these results, Diffusion-Based Impedance Learning establishes a general interface between learning and model-based control. It extends the role of generative models from visuomotor trajectory planning to physical interaction, demonstrating that contact-rich manipulation can be learned from relatively small datasets while maintaining real-time control performance.
Looking ahead, we aim to extend this framework from single-arm to dual-arm setups. While this work focuses on impedance learning and used a fixed ZFT, Diffusion-Based Impedance Learning can naturally be embedded in hierarchical architectures that couple high-level visuomotor motion policies with low-level impedance regulation.
Appendix A ZFT-based teleoperation with Apple Vision Pro
@@
Figure 12: Pose tracking and visualization with Apple Vision Pro. Left: operator performing teleoperation. Right: virtual reconstruction of wrist, hand, and head poses in a 3D scene. [55]
@@
Pose data were transferred via shared memory between the Python-based VisionProTeleop module and the C++ robot controller. In this way, the Apple Vision Pro provided a seamless method to generate sZFT ground truth data through markerless and physically unconstrained teleoperation. The implementation of the teleoperation interface is available in our GitHub repository [^16] https://GitHub.com/diffusion-imp-anon/DiffusionBasedImpedanceLearning.
Appendix B Peg-In-Hole experiments
This appendix complements the main results by reporting stiffness adaptation for the circular and square pegs. The corresponding results for the star peg are already presented in the main text (Figure 11). The experimental setup, controllers, and parameters matched the main text: baseline stiffness $K_{t,i,\text{max}}=$800\text{\,}\mathrm{N}\mathrm{/}\mathrm{m}$\$ and $K_{r,i,\text{max}}=$150\text{\,}\mathrm{N}\mathrm{m}\mathrm{/}\mathrm{r}\mathrm{a}\mathrm{d}$\$, damping as in Appendix II-B1, and force/moment thresholds of $1\text{\,}\mathrm{N}$ and $1\text{\,}\mathrm{N}\mathrm{m}$. All quantities are expressed in the robot base frame.
@@ Appendix B Peg-In-Hole experiments
Figure 13: Cylindrical peg insertion with stiffness adaptation. External wrenches, commanded translational and rotational stiffness, and relative motion factors are shown in the robot base frame. Base-frame color coding is given in Figure 9.
Figure 14: Square peg insertion with stiffness adaptation. External wrenches, commanded translational and rotational stiffness, and relative motion factors are shown in the robot base frame. Base-frame color coding is given in Figure 9.